In a sample of 800 adults, 228 said that they liked sugar cereals. Three adults are selected at random without replacement. Find the probability that all three like sugar cereals.
Answer #1

ans) Here total numbers of adults are 800

adults who likes  sugar cereals is 228

if we select the three adults (without replacement) who like sugar cereals then

P1(adults likes sugar cereals) = 228/800

now there are 227 remaining adults who like sugar cereals out of 799 adults then

P2((adults likes sugar cereals) =227/799

similiarly,

there are 226 remaining adults who like sugar cereals out of 798 adults then

P3((adults likes sugar cereals) =226/798

probability that all three like sugar cereals = P1*P2*P3

=228*227*226/800*799*798

=11696856/510081600

=0.0229 =2.29%